# Slender Trees and the Approximation Property

### v1

@@ 5. Controlling the Slenderness at $\kappa$
Clearly $\mathbb{P}$ is of size $\kappa$ and $\kappa$-cc. Let $e$ be a $(\kappa,\lambda)$-list. Let $\Theta$ be large. Choose any $\lambda$-ineffability witness $M$ for $\kappa$ with respect to $e$ as in Lemma 4.3. If $G$ is $\mathbb{M}_{0}(\tau,\mu,\kappa)$-generic over $V$, $\pi[G\cap M]$ is $\pi(\mathbb{M}_{0}(\tau,\mu,\kappa))=\mathbb{M}_{0}(\tau,\mu,M\cap\kappa)$-generic over $V$. By Lemma 5.6, $V[G]$ is an extension of $V[\pi[G\cap M]]$ by forcing with $\mathbb{M}_{0}(G,\tau,\mu,\kappa\smallsetminus(M\cap\kappa))$, which has the ${<}\,\tau^{+}$-approximation property. ∎
@@
Let $(x_{\alpha})_{\alpha<\theta}$ enumerate different subsets of $2^{\delta}$. Let $e$ be the following $(\theta,\theta)$-list: $e(a)=x_{a}$ if $a$ is an ordinal and $\emptyset$ otherwise.
@@
Let $\Theta$ be large and $C$ be the club of all $M\in[H(\Theta)]^{<\theta}$ such that $[\delta]^{<\delta}\subseteq M$ (here we use $2^{<\delta}<\theta$). If $M\in C$ and $x\in[\theta]^{<\delta}$, $e(M\cap\theta)\cap x\in[\delta]^{<\delta}$ (since $e(M\cap\theta)$ is either a subset of $\delta$ or empty) and thus in $M$. ∎
@@
Otherwise there is $b$ such that $S:=\{a\in[\theta]^{<\theta}\;|\;e(a)=b\cap a\}$ is stationary in $[\theta]^{<\theta}$. In particular $S\cap\theta$ is stationary in $\theta$. However, since $e(a)$ is a subset of $\delta$ for every $a$, $e$ is constant on $(S\cap\theta)\smallsetminus\delta$ which is an obvious contradiction. ∎
@@
Lemma 5.9 follows because $\tau^{<\tau}=\tau<\kappa$ in $V$, which is preserved into $V[G]$ by the ${<}\,\tau$-directed closure of $\mathbb{M}_{0}(\tau,\mu,\kappa)$. We obtain another Corollary:
@@ 6. $\operatorname{ISP}$ and Cardinal Arithmetic
For the rest of this section, fix regular cardinals $\tau<\mu<\kappa\leq\lambda$ such that $\tau^{<\tau}=\tau$, $\mu^{<\mu}=\mu$ and $\kappa$ is inaccessible. Also use the same definition for $\operatorname{Add}(\tau,\gamma)$ as in the previous section. For $\mathbb{M}_{1}$, we can show directly that the “quotient ordering“ has the correct approximation property (of course building on the results for $\mathbb{M}_{0}$).
@@
Let $\nu\in(\mu,\kappa]$ be an inaccessible cardinal and $\gamma$ an ordinal. Let $G^{\prime}\times H^{\prime}$ be $\mathbb{M}_{1}(\tau,\mu,\nu,\gamma)$-generic. In $V[G^{\prime}\times H^{\prime}]$, $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus\nu)\times\operatorname{Add}(\mu,\lambda\smallsetminus\gamma)$ has the ${<}\,\nu$-approximation property.
@@
$\operatorname{Add}(\mu,\lambda\smallsetminus\gamma)$ is $\nu$-Knaster in $V[G^{\prime}\times H^{\prime}]$ and thus has the ${<}\,\nu$-approximation property. Let $H^{\prime\prime}$ be $\operatorname{Add}(\mu,\lambda\smallsetminus a)$-generic over $V[G^{\prime}\times H^{\prime}]$. Then $V[G^{\prime}\times H^{\prime}][H^{\prime\prime}]$ is equal to $V[G^{\prime}][H]$, where $H$ is $\operatorname{Add}(\mu,\lambda)^{V}$-generic (which is ${<}\,\mu$-distributive in $V[G^{\prime}]$). In $V[G^{\prime}]$, $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus\nu)$ is iteration-like and has $\operatorname{Add}(\tau,\kappa\smallsetminus\nu)$ as its base ordering as well as a ${<}\,\mu$-closed term ordering. Thus, in $V[G^{\prime}][H]$, the ordering is still iteration-like (as this property is absolute), the base ordering is still $\tau^{+}$-Knaster and the term ordering is still ${<}\,\mu$-closed. Hence $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus\nu)$ has the ${<}\,\tau^{+}$-approximation property in $V[G^{\prime}][H]=V[G^{\prime}\times H^{\prime}][H^{\prime\prime}]$. Now assume there is $f\in V[G\times H]$ such that $f\cap z\in V[G^{\prime}\times H^{\prime}]$ for every $z\in[V[G^{\prime}\times H^{\prime}]]^{<\nu}\cap V[G^{\prime}\times H^{\prime}]$. Let $z\in V[G^{\prime}\times H]$ have size ${<}\,\nu$. Because $\operatorname{Add}(\mu,\lambda)$ is $\nu$-Knaster there is $y\in V[G^{\prime}\times H^{\prime}]$ with $z\subseteq y$ and $|y|<\nu$. Hence $f\cap y\in V[G^{\prime}\times H^{\prime}]$ and $f\cap z=(f\cap y)\cap z\in V[G^{\prime}\times H]$. As $z$ was arbitrary and $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus\nu)$ has the ${<},\tau^{+}$-approximation property in $V[G^{\prime}\times H^{\prime}]$, $f\in V[G^{\prime}\times H]$. Now since $\operatorname{Add}(\mu,\lambda\smallsetminus\gamma)$ has the ${<}\,\nu$-approximation property in $V[G^{\prime}\times H^{\prime}]$, $f\in V[G^{\prime}\times H^{\prime}]$. ∎
@@
Again, $\mathbb{P}$ has size $\leq\lambda_{0}$ and is $\kappa$-cc.. Let $e$ be any $(\kappa,\lambda)$-list, $\Theta$ large and $x\in[H(\Theta)]^{<\kappa}$. Choose any $\lambda$-ineffability witness $M$ for $\kappa$ with respect to $e$ as in Lemma 4.3 with $\tau^{+}\in M$. We have $\pi(\mathbb{M}_{1}(\tau,\mu,\kappa,\lambda))=\mathbb{M}_{0}(\tau,\mu,M\cap\kappa)\times\operatorname{Add}(\mu,\pi(\lambda))$. If $G$ is $\mathbb{M}_{1}(\tau,\mu,\kappa,\lambda)$-generic, then $\pi[G\cap M]$ is $\mathbb{M}_{0}(\tau,\mu,M\cap\kappa)\times\operatorname{Add}(\mu,\pi(\lambda))$-generic by its cc.. $V[G]$ is an extension of $V[\pi[G\cap M]]$ by $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus(M\cap\kappa))\times\operatorname{Add}(\mu,\lambda\smallsetminus\pi(\lambda))$ which has the ${<}\,M\cap\kappa$-approximation property in $V[\pi[G\cap M]]$ by Lemma 6.2. ∎
@@
As before, let $f(a)$ be the $a$th Cohen subset of $\mu$ added by $\mathbb{M}_{1}(\tau,\mu,\kappa,\lambda)$ if $a$ is an ordinal and $\emptyset$ otherwise. Because every ${<}\,\mu$-sized segment of $f(a)$ is in $([\mu]^{<\mu})^{V}$ which has size $\mu<\kappa$, we see that $f$ is $\mu$-slender. However, by previous arguments $f$ cannot have an ineffable branch. ∎
@@
For the rest of this section, we drop the assumption that $\mu^{<\mu}=\mu$. We have a very similar Lemma to before (albeit with a stronger approximation property):
@@
Let $\nu\in(\mu,\kappa)$ be inaccessible and let $G^{\prime}\times H^{\prime}$ be $\mathbb{M}_{2}(\tau,\mu,\nu,\gamma)$-generic. In $V[G^{\prime}\times H^{\prime}]$, $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus\nu)\times\operatorname{Add}(\tau,\lambda\smallsetminus\gamma)$ has the ${<}\,\tau^{+}$-approximation property.
@@
In $V[G^{\prime}\times H^{\prime}]$, $\operatorname{Add}(\tau,\lambda\smallsetminus a)$ is still $(2^{<\tau})^{V[G^{\prime}\times H^{\prime}]}=(\tau^{+})^{V[G^{\prime}\times H^{\prime}]}=(\tau^{+})^{V}$-cc., so $\operatorname{Add}(\tau,\lambda\smallsetminus a)$ has the ${<}\,\tau^{+}$-approximation property in $V[G^{\prime}\times H^{\prime}]$. Let $H^{\prime\prime}$ be $\operatorname{Add}(\tau,\lambda\smallsetminus a)$ be $\operatorname{Add}(\tau,\lambda\smallsetminus a)$-generic and $H$ the $\operatorname{Add}(\tau,\lambda)$-generic filter induced by $H^{\prime}$ and $H^{\prime\prime}$. In $V[G^{\prime}]$, $\mathbb{M}_{0}(G^{\prime},\tau,\tau^{+},\kappa\smallsetminus\nu)$ has a ${<}\,\tau^{+}$-Knaster base ordering and a ${<}\,\tau^{+}$-closed term ordering. In $V[G^{\prime}\times H]$, the base ordering is still ${<}\,\tau^{+}$-Knaster and the term ordering is at least ${<}\,\tau^{+}$-strongly distributive, because $V[G^{\prime}\times H]$ is an extension of $V[G^{\prime}]$ by a $\tau^{+}$-Knaster forcing. Being iteration-like is absolute and thus $\mathbb{M}_{0}(G^{\prime},\tau,\tau^{+},\kappa\smallsetminus\nu)$ has the ${<}\,\tau^{+}$-approximation property in $V[G^{\prime}\times H]$. Now proceed as in Lemma 6.2. ∎
@@
7. Making $\operatorname{ISP}$ indestructible under directed closed forcing
It is a well-known result by Laver that if $\kappa$ is a supercompact cardinal, there is a forcing which leaves $\kappa$ supercompact and moreover makes the supercompactness of $\kappa$ indestructible under ${<}\,\kappa$-directed closed forcing. Unger showed in [14] that it is possible to force the tree property at $\omega_{2}$ such that it is indestructible under ${<}\,\omega_{2}$-directed closed forcing. We will adapt his arguments to show that it is consistent that $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$ holds (for $\kappa$ a successor of a regular cardinal), where $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$ means that $\operatorname{ISP}(\tau^{+},\kappa,\lambda)$ holds for every $\lambda\geq\kappa$, and is indestructible under ${<}\,\kappa$-directed closed forcing. Our forcing will be a guessing variant of Mitchell forcing, modified to collapse cardinals in a “non-fresh“ way to ensure the approximation property. We will need the concept of a Laver function:

### v2

@@ 5. Controlling the Slenderness at $\kappa$
Clearly $\mathbb{M}_{0}(\tau,\mu,\kappa)$ is of size $\kappa$ and $\kappa$-cc. Let $e$ be a $(\kappa,\lambda)$-list. Let $\Theta$ be large. Choose any $\lambda$-ineffability witness $M$ for $\kappa$ with respect to $e$ as in Lemma 4.3. If $G$ is $\mathbb{M}_{0}(\tau,\mu,\kappa)$-generic over $V$, $\pi[G\cap M]$ is $\pi(\mathbb{M}_{0}(\tau,\mu,\kappa))=\mathbb{M}_{0}(\tau,\mu,M\cap\kappa)$-generic over $V$. By Lemma 5.6, $V[G]$ is an extension of $V[\pi[G\cap M]]$ by forcing with $\mathbb{M}_{0}(G,\tau,\mu,\kappa\smallsetminus(M\cap\kappa))$, which has the $\tau^{+}-$approximation property by Corollary 5.5. So by Theorem 4.8, $\mathbb{M}_{0}(\tau,\mu,\kappa)$ forces $\operatorname{ISP}(\tau^{+},\kappa,\lambda)$. ∎
@@
Let $(x_{\alpha})_{\alpha<\theta}$ enumerate different subsets of $\delta$. Let $e$ be the following $(\theta,\theta)$-list: $e(a)=x_{a}$ if $a\geq\delta$ is an ordinal and $\emptyset$ otherwise.
@@
Let $\Theta$ be large and $C$ be the club of all $M\in[H(\Theta)]^{<\theta}$ such that $\delta]^{<\delta}\subseteq M$ (here we use $2^{<\delta}<\theta$). If $M\in C$ and $x\in[\theta]^{<\delta}$, $e(M\cap\theta)\cap x\in[\delta]^{<\delta}$ (since $e(M\cap\theta)$ is either a subset of $\delta$ or empty) and thus in $M$. ∎
@@
Otherwise there is $b$ such that $S:=\{a\in[\theta]^{<\theta}\;|\;e(a)=b\cap a\}$ is stationary in $\theta]^{<\theta}$. In particular $S\cap\theta$ is stationary in $\theta$. However, since $e(a)$ is a subset of $\delta$ for every $a$, $e$ is constant on $(S\cap\theta)\smallsetminus\delta$ which is an obvious contradiction. ∎
@@
Lemma 5.9 follows because $\tau^{<\tau}=\tau<\kappa$ in $V$, which is preserved into $V[G]$ by the $\tau-$directed closure of $\mathbb{M}_{0}(\tau,\mu,\kappa)$. We obtain another Corollary:
@@ 6. $\operatorname{ISP}$ and Cardinal Arithmetic
For the rest of this section, fix regular cardinals $\tau<\mu<\kappa\leq\lambda$ such that $\tau^{<\tau}=\tau$, $\mu^{<\mu}=\mu$ and $\kappa$ is inaccessible. Also use the same definition for $\operatorname{Add}(\tau,\gamma)$ as in the previous section.
For $\mathbb{M}_{1}$, we can show directly that the “quotient ordering“ has the correct approximation property (of course building on the results for $\mathbb{M}_{0}$).
@@
Let $\nu\in(\mu,\kappa]$ be an inaccessible cardinal and $\gamma$ an ordinal. Let $G^{\prime}\times H^{\prime}$ be $\mathbb{M}_{1}(\tau,\mu,\nu,\gamma)$-generic. In $V[G^{\prime}\times H^{\prime}]$, $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus\nu)\times\operatorname{Add}(\mu,\lambda\smallsetminus\gamma)$ has the $\nu-$approximation property.
@@
$\operatorname{Add}(\mu,\lambda\smallsetminus\gamma)$ is $\nu$-Knaster in $V[G^{\prime}\times H^{\prime}]$ and thus has the $\nu-$approximation property. Let $H^{\prime\prime}$ be $\operatorname{Add}(\mu,\lambda\smallsetminus\gamma)$-generic over $V[G^{\prime}\times H^{\prime}]$. Then $V[G^{\prime}\times H^{\prime}][H^{\prime\prime}]$ is equal to $V[G^{\prime}][H]$, where $H$ is $\operatorname{Add}(\mu,\lambda)^{V}$-generic over $V[G^{\prime}]$. We note that $\operatorname{Add}(\mu,\lambda)^{V}$ is $\mu-$distributive in $V[G^{\prime}]$: By Lemma 5.2 and Lemma 3.7 $V[G^{\prime}]$ is contained in an extension by the product of a $\mu-$closed and a $\tau^{+}$-Knaster forcing. In that extension, $\operatorname{Add}(\mu,\lambda)^{V}$ is clearly $\mu-$distributive by Easton’s Lemma and thus the same statement holds in the smaller model $V[G^{\prime}]$.
In $V[G^{\prime}]$, $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus\nu)$ is iteration-like and has $\operatorname{Add}(\tau,\kappa\smallsetminus\nu)$ as its base ordering as well as a $\mu-$closed term ordering (by Lemma 5.4 and its proof). Thus, in $V[G^{\prime}][H]$, the ordering is still iteration-like (as this property is absolute), the base ordering is still $\tau^{+}$-Knaster and the term ordering is still $\mu-$closed. Hence $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus\nu)$ has the $\tau^{+}-$approximation property in $V[G^{\prime}][H]=V[G^{\prime}\times H^{\prime}][H^{\prime\prime}]$.
Now assume there is $f\in V[G\times H]$ such that $f\cap z\in V[G^{\prime}\times H^{\prime}]$ for every $z\in[V[G^{\prime}\times H^{\prime}]]^{<\nu}\cap V[G^{\prime}\times H^{\prime}]$. Let $z\in V[G^{\prime}\times H]$ have size $\nu$. Because $\operatorname{Add}(\mu,\lambda)$ is $\nu$-Knaster there is $y\in V[G^{\prime}\times H^{\prime}]$ with $z\subseteq y$ and $|y|<\nu$. Hence $f\cap y\in V[G^{\prime}\times H^{\prime}]$ and $f\cap z=(f\cap y)\cap z\in V[G^{\prime}\times H]$. As $z$ was arbitrary and $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus\nu)$ has the $\tau^{+}-$approximation property in $V[G^{\prime}\times H^{\prime}]$, $f\in V[G^{\prime}\times H]$. Now since $\operatorname{Add}(\mu,\lambda\smallsetminus\gamma)$ has the $\nu-$approximation property in $V[G^{\prime}\times H^{\prime}]$, $f\in V[G^{\prime}\times H^{\prime}]$. ∎
@@
Again, $\mathbb{M}_{1}(\tau,\mu,\kappa,\lambda)$ has size $\leq\lambda_{0}$ and is $\kappa$-cc since it is the product of two $\kappa$-Knaster posets (this is clear for $\operatorname{Add}(\mu,\lambda)$ and was shown in Lemma 5.2 for $\mathbb{M}_{0}(\tau,\mu,\kappa)$). Let $e$ be any $(\kappa,\lambda)$-list, $\Theta$ large and $x\in[H(\Theta)]^{<\kappa}$. Choose any $\lambda$-ineffability witness $M$ for $\kappa$ with respect to $e$ as in Lemma 4.3 with $\tau^{+}\in M$. We have $\pi(\mathbb{M}_{1}(\tau,\mu,\kappa,\lambda))=\mathbb{M}_{0}(\tau,\mu,M\cap\kappa)\times\operatorname{Add}(\mu,\pi(\lambda))$. If $G$ is $\mathbb{M}_{1}(\tau,\mu,\kappa,\lambda)$-generic, then $\pi[G\cap M]$ is $\mathbb{M}_{0}(\tau,\mu,M\cap\kappa)\times\operatorname{Add}(\mu,\pi(\lambda))$-generic over $V$ by its $M\cap\kappa$-cc (again, it is the product of two $M\cap\kappa$-Knaster posets). $V[G]$ is an extension of $V[\pi[G\cap M]]$ by $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus(M\cap\kappa))\times\operatorname{Add}(\mu,\lambda\smallsetminus\pi(\lambda))$ which has the $M\cap\kappa-$approximation property in $V[\pi[G\cap M]]$ by Lemma 6.2. So we can once again apply Theorem 4.8. ∎
@@
As before, let $f(a)$ be the $a$th Cohen subset of $\mu$ added by $\mathbb{M}_{1}(\tau,\mu,\kappa,\lambda)$ if $a\geq\mu$ is an ordinal and $\emptyset$ otherwise. Because every $\mu-$sized segment of $f(a)$ is in $([\mu]^{<\mu})^{V}$ which has size $\mu<\kappa$, we see that $f$ is $\mu$-slender. However, by previous arguments $f$ cannot have an ineffable branch. ∎
@@
For the rest of this section, we drop the assumption that $\mu^{<\mu}=\mu$.
We have a very similar Lemma to before (albeit with a stronger approximation property):
@@
Let $\nu\in(\mu,\kappa)$ be inaccessible and let $G^{\prime}\times H^{\prime}$ be $\mathbb{M}_{2}(\tau,\mu,\nu,\gamma)$-generic. In $V[G^{\prime}\times H^{\prime}]$, $\mathbb{M}_{0}(G^{\prime},\tau,\mu,\kappa\smallsetminus\nu)\times\operatorname{Add}(\tau,\lambda\smallsetminus\gamma)$ has the $\tau^{+}-$approximation property.
@@
In $V[G^{\prime}\times H^{\prime}]$, $\operatorname{Add}(\tau,\lambda\smallsetminus\gamma)$ is still $(2^{<\tau})^{V[G^{\prime}\times H^{\prime}]}=(\tau^{+})^{V[G^{\prime}\times H^{\prime}]}=(\tau^{+})^{V}$-cc., so $\operatorname{Add}(\tau,\lambda\smallsetminus\gamma)$ has the $\tau^{+}-$approximation property in $V[G^{\prime}\times H^{\prime}]$. Let $H^{\prime\prime}$ be $\operatorname{Add}(\tau,\lambda\smallsetminus\gamma)$ be $\operatorname{Add}(\tau,\lambda\smallsetminus\gamma)$-generic and $H$ the $\operatorname{Add}(\tau,\lambda)$-generic filter induced by $H^{\prime}$ and $H^{\prime\prime}$. In $V[G^{\prime}]$, $\mathbb{M}_{0}(G^{\prime},\tau,\tau^{+},\kappa\smallsetminus\nu)$ has a $\tau^{+}-$Knaster base ordering and a $\tau^{+}-$closed term ordering. In $V[G^{\prime}\times H]$, the base ordering is still $\tau^{+}-$Knaster and the term ordering is at least $\tau^{+}-$strongly distributive, because $V[G^{\prime}\times H]$ is an extension of $V[G^{\prime}]$ by a $\tau^{+}$-Knaster forcing. Being iteration-like is absolute and thus $\mathbb{M}_{0}(G^{\prime},\tau,\tau^{+},\kappa\smallsetminus\nu)$ has the $\tau^{+}-$approximation property in $V[G^{\prime}\times H]$. Now proceed as in Lemma 6.2. ∎
@@
7. Indestructibility of $\operatorname{ISP}$
It is a well-known result by Laver that if $\kappa$ is a supercompact cardinal, there is a forcing which leaves $\kappa$ supercompact and moreover makes the supercompactness of $\kappa$ indestructible under $\kappa-$directed closed forcing (see [15]). It is also known that if the proper forcing axiom $\operatorname{PFA}$ holds, the tree property at $\aleph_{2}$ is indestructible by $\aleph_{2}-$closed forcing (see [13]). In [21], Unger devised a guessing variant of Mitchell forcing to present a more adaptable method for obtaining the indestructibility of the tree property under directed-closed forcings (in particular, this method can be carried out at cardinals above $\aleph_{2}$). We will adapt his arguments to show that it is consistent that $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$ holds (for $\kappa$ a successor of a regular cardinal), where $\operatorname{ISP}(\tau^{+},\kappa,\geq\kappa)$ means that $\operatorname{ISP}(\tau^{+},\kappa,\lambda)$ holds for every $\lambda\geq\kappa$, and is indestructible under $\kappa-$directed closed forcing.
Our forcing will be a guessing variant of Mitchell forcing, modified to collapse cardinals in a “non-fresh“ way to ensure the approximation property.
We will need the concept of a Laver function: